OpenStudy (cutegirl1000):

Which step should be completed first to solve the equation ? A. Add 2 to both sides. B. Subtract 4 from both sides. C. Divide both sides by 3. D. Add 4 to both sides.

OpenStudy (cutegirl1000):

can you explain???

OpenStudy (welshfella):

- please dont give direct answers

OpenStudy (cutegirl1000):

yeah I want to understand this topic

OpenStudy (elijahrich1113):

So if you want to get z by itself you have to subtract 4 from both sides

OpenStudy (elijahrich1113):

Do you know what to do next?

OpenStudy (cutegirl1000):

so -6 = 3z -2 = z???

OpenStudy (elijahrich1113):

yes good job :)

OpenStudy (cutegirl1000):

thanks can you help me with 4 more?

OpenStudy (elijahrich1113):

sure

OpenStudy (cutegirl1000):

thanks

OpenStudy (elijahrich1113):

just post them in the comments here

OpenStudy (cutegirl1000):

What is the solution to the equation ? x = https://static.k12.com/bank_packages/files/media/mathml_5d691dfadf7468d726f272cfdbd9060f09c2247b_1.gif

OpenStudy (elijahrich1113):

do you know the distributive property?

OpenStudy (cutegirl1000):

one sec I am trying to figure out an answer so I can see if i'm correct

OpenStudy (cutegirl1000):

yeah

OpenStudy (cutegirl1000):

oh wait... one sec again

OpenStudy (elijahrich1113):

what are the steps you took cause I got a diff. answer

OpenStudy (cutegirl1000):

I got x =5

OpenStudy (elijahrich1113):

yes :)

OpenStudy (cutegirl1000):

What is the solution to the equation ? m = https://static.k12.com/bank_packages/files/media/mathml_34efec2888329d378c42ec5edf2f4e98683a8116_1.gif

OpenStudy (cutegirl1000):

i think it is...

OpenStudy (cutegirl1000):

... m = -4

OpenStudy (elijahrich1113):

yes :)

OpenStudy (elijahrich1113):

so yeah that is pretty much the hang of it

OpenStudy (cutegirl1000):

Which answer best describes the solution to the equation 5z + 4 = 4 + 5z? A. one solution B. no solution C. an infinite number of solutions D. z = 20

OpenStudy (cutegirl1000):

I think it is...

OpenStudy (cutegirl1000):

is it c?

OpenStudy (elijahrich1113):

yes

OpenStudy (cutegirl1000):

one more

OpenStudy (elijahrich1113):

ok

OpenStudy (elijahrich1113):

are you only supposed to answer 3 of them or is there more than one choice for an answer?

OpenStudy (cutegirl1000):

one sec I am working on it

OpenStudy (elijahrich1113):

I have the answers to check

OpenStudy (cutegirl1000):

it is B, C, A, A

OpenStudy (elijahrich1113):

I got B, C, B, A

OpenStudy (cutegirl1000):

No 3. you first multiply by 4 then you get x + 3 = 3x + 3

OpenStudy (cutegirl1000):

then you move x to the other side so you do 3 = 2x + 3 0 = 2x x = 0 One solution

OpenStudy (elijahrich1113):

I multiplied the 4 to the 3/4 which gives you 3. Then I subtracted 3/4x from x to get 1/4x. Then you subtract the 3 which gives you zero and then 1/4 does not equal 0?

OpenStudy (elijahrich1113):

But good luck I gtg and I hope k12 isnt hard cause im doing k12 next year!!!

OpenStudy (cutegirl1000):

wait... you multiply 3/4 x by 4 and get 3x

OpenStudy (cutegirl1000):

@ElijahRich1113

OpenStudy (cutegirl1000):

I got 100% so thanks for all your help the only thing is that I was right about the third one being A not B

OpenStudy (elijahrich1113):

Oh ok good 👍🏽